# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ G.1 Main Proof of Theorem 3
Theorem G.1 (Lower bound on adaptive PL model parameter estimation).
@@
Proof of Theorem G.1.
The analysis of this result closely follows the proof of Theorem 3 in Shamir [75]. The key distinction lies in the input structure: our setting involves a set of feature vectors, while theirs is restricted to a single feature vector.
@@
$\geq\mu^{2}\cdot\mathbb{E}_{\bm{\theta}\sim\operatorname{Unif}(\Theta)}\mathbb{E}_{\bm{\theta}}\left[\sum_{i=1}^{d}\mathbb{I}\left\{\bm{\theta}_{i}\widehat{\bm{\theta}}_{i}<0\right\}\right].$ (G.1)
As in Shamir [75], we assume that the query strategy is deterministic conditioned on the past: that is, $S_{t}$ is a deterministic function of the previous queries and observations, i.e., $S_{1},\sigma_{1},\dots,S_{t-1},\sigma_{t-1}$. This assumption is made without loss of generality, since any randomized querying strategy can be viewed as a distribution over deterministic strategies. Therefore, a lower bound that holds uniformly for all deterministic strategies also applies to any randomized strategy. Then, we use the following lemma.
Lemma G.1 (Lemma 4 of Shamir 75).
@@
The proof is deferred to Appendix G.2.1.
By applying Lemma G.2, we have
@@
$\geq\frac{d}{2}\left(1-\sqrt{\frac{2K^{2}T\mu^{2}}{d}}\right).$ (G.2)
Combining Equation (G.1) and (G.2), we prove the first inequality of Theorem G.1. The second inequality directly follows by choosing $\mu=\sqrt{d/(8K^{2}T)}$. ∎
@@
The structure of our proof is similar to that of Theorem 2 in Wagenmaker et al. [81]. However, while they consider the linear bandit setting, we focus on the Plackett–Luce (PL) bandit setting.
We adopt the same instance construction as in Theorem G.1, where $\Phi={\mathcal{S}}^{d-1}$ and $\Theta=\{-\mu,\mu\}^{d}$. Define $\phi^{\star}(\bm{\theta})=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi_{a}^{\top}\bm{\theta}$. Then, since $\phi^{\star}(\bm{\theta})\in\Phi$ and $\bm{\theta}\in\Phi$, it is clear that
$\phi^{\star}(\bm{\theta})=\bm{\theta}/\|\bm{\theta}\|_{2}=\bm{\theta}/(\sqrt{d}\mu),\quad\text{and }\,\phi^{\star}(\bm{\theta})^{\top}\bm{\theta}=\sqrt{d}\mu.$ (G.3)
@@
$\mathbb{E}_{\phi\sim\pi}\left[\phi^{\top}\bm{\theta}\right]=\big(\underbrace{\mathbb{E}_{\phi\sim\pi}\left[\phi\right]}_{=:\phi_{\pi}}\big)^{\top}\bm{\theta}\geq\phi^{\star}(\bm{\theta})^{\top}\bm{\theta}-\epsilon=\sqrt{d}\mu-\epsilon.$ (G.4)
@@
$\Longleftrightarrow\,\,$ $\bm{\theta}^{\star}\Delta\leq-\frac{\sqrt{d}\mu}{2}\|\Delta\|_{2}^{2}.$ (Eqn. (G.3))
@@
$-\epsilon\leq-\frac{\sqrt{d}\mu}{2}\|\Delta\|_{2}^{2}.$ (Eqn. (G.4))
@@
$\big\|\widehat{\bm{\theta}}-\bm{\theta}\big\|_{2}=\big\|\sqrt{d}\mu(\widehat{\phi}-\Delta^{\prime})-\sqrt{d}\mu(\widehat{\phi}-\Delta)\big\|_{2}\leq 2\sqrt{d}\mu\sqrt{\frac{2\epsilon}{\sqrt{d}\mu}}=\sqrt{8\sqrt{d}\mu\epsilon}.$ (G.5)
@@
$\leq 8\sqrt{d}\mu\epsilon+\mathbb{E}_{\bm{\theta}}\left[\big\|\widehat{\bm{\theta}}-\bm{\theta}\big\|_{2}^{2}\cdot\mathbb{I}\{\mathcal{E}^{c}\}\right]$ (Eqn. (G.5))
@@
On the other hand, by Theorem G.1, there exists a parameter $\bm{\theta}\in\Theta$ such that, if we collect $T$ samples and set $\mu=\sqrt{d/(8K^{2}T)}$, then the following lower bound holds:
@@
G.2 Proof of Lemmas for Theorem 3
G.2.1 Proof of Lemma G.2
Proof of Lemma G.2.
@@ G.2 Proof of Lemmas for Theorem 3
$D_{\text{KL}}\big(\mathbb{P}_{\bm{\theta}}(\cdot|S)\|\mathbb{P}_{\bm{\theta}^{\prime}}(\cdot|S)\big)=\mathbb{E}_{\sigma\sim\mathbb{P}_{\bm{\theta}}(\cdot|S)}\left[\sum_{j=1}^{|S|}\left(\phi_{\sigma_{j}}^{\top}\left(\bm{\theta}-\bm{\theta}^{\prime}\right)-\log\frac{\sum_{k=j}^{|S|}e^{\phi_{\sigma_{k}}^{\top}\bm{\theta}}}{\sum_{k=j}^{|S|}e^{\phi_{\sigma_{k}}^{\top}\bm{\theta}^{\prime}}}\right)\right].$ (G.6)
@@
$\leq\sum_{k=j}^{|S|}p_{k}(\bm{\theta})\phi_{\sigma_{k}}^{\top}\left(\bm{\theta}^{\prime}-\bm{\theta}\right)+\frac{1}{2}\sum_{a\in S}\left(\phi_{a}^{\top}(\bm{\theta}^{\prime}-\bm{\theta})\right)^{2},$ (G.7)
@@
Plugging Equation (G.7) into Equation (G.6), we get
@@
Appendix H Additional Discussions
In this section, we provide additional discussion of our approach. In Subsection H.1, we propose a more efficient assortment selection rule than Equation (9), by using an arbitrary reference action $\bar{a}_{t}\in\mathcal{A}$ instead of selecting the one that maximizes average uncertainty. In Subsection H.2, we show that under a sufficient feature diversity condition, selecting $S_{t}$ uniformly at random can still achieve a comparable suboptimality gap. Finally, in Subsection H.3, we extend our approach to the active learning setting, as studied in [19].
H.1 Arbitrary Reference Action for More Efficient Assortment Selection
As described in the main paper, the reference action $\bar{a}_{t}$ is selected to maximize the average uncertainty across the subset $S_{t}$, according to Equation (9). This selection incurs a computational cost of $\tilde{\mathcal{O}}(N^{2}K)$ (see Remark 2).
However, in this subsection, we show that $\bar{a}_{t}$ can, in fact, be selected arbitrarily—i.e., any $\bar{a}_{t}\in\mathcal{A}$ is valid. Specifically, we modify our assortment selection rule as follows:
$S_{t}=\mathop{\mathrm{argmax}}_{\begin{subarray}{c}S\in{\mathcal{S}}\\ \bar{a}_{t}\in S\end{subarray}}\frac{1}{|S|}\sum_{a\in S}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}},\quad\text{for any }\bar{a}_{t}\in\mathcal{A}.$ (H.1)
This results in only a constant-factor increase (specifically, by a factor of 2) in the suboptimality gap, while reducing the computational cost to $\tilde{\mathcal{O}}(NK)$, as it removes the need to enumerate over all possible reference actions.
To show this explicitly, we return to Equation (D.6). Let $\bar{a}_{t}$ be an arbitrary action in $\mathcal{A}$ (e.g., selected uniformly at random). Then, we have
$\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\pm\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}$
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}+\left\|\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}\right)$
$=\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\Bigg(\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}+\sum_{a^{\prime}\in\widehat{S}_{t}}\left\|\phi\left(x_{t},a^{\prime}\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}\Bigg)$ (Let $S^{\star}_{t}:=\{\pi^{\star}(x_{t}),\bar{a}_{t}\}$ and $\widehat{S}_{t}:=\{\widehat{\pi}_{T}(x_{t}),\bar{a}_{t}\}$ )
$\leq\frac{4\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}}.$ ($S_{t}$ selection rule, Eqn. (H.1))
The remaining steps of the proof follow exactly as in the proof of Theorem 1 (or Theorem 2).
H.2 Suboptimality Gap Under Sufficient Diversity Condition
So far, we have considered the general case where the feature vectors $\phi$ are not required to be diverse, and as a result, the induced matrix $H_{t}-\lambda\mathbf{I}_{d}$ (or $\Lambda_{t}-\lambda\mathbf{I}_{d}$) may be singular. In this subsection, we discuss the case where the following diversity assumption holds:
Assumption H.1 (Diverse features).
For any $S\in{\mathcal{S}}$ and $a^{\prime}\in S$, there exists a constant $\lambda_{0}>0$ such that $\lambda_{\min}\left(\mathbb{E}_{x\sim\rho}\left[\frac{1}{|S|}\sum_{a\in S}(\phi(x,a)-\phi(x,a^{\prime}))(\phi(x,a)-\phi(x,a^{\prime}))^{\top}\right]\right)\geq\lambda_{0}$.
Under this condition, it is sufficient to randomly select exactly $K$ actions, rather than solving the optimization problem in Equation (9) to construct the assortment. Specifically, we can select $S_{t}$ as:
$S_{t}\sim\text{Unif}\left(\{S\subseteq\mathcal{A}:|S|=K\}\right),\quad\forall t\in[T].$ (H.2)
Theorem H.1 (Suboptimality Gap of Random Assortment Selection Under Diversity).
Let ${\mathcal{T}}^{w}:=\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\right\}$, where $a_{t}^{\prime}\in S_{t}$ is an arbitrary action selected from the assortment $S_{t}$. Suppose $T=\Omega\big(\log(dT)/\lambda_{0}\big)$ and $T>|{\mathcal{T}}^{w}|$. Then, under the same setting as Theorem 1 and Assumption H.1, if $S_{t}$ is randomly selected according to Equation (H.2), then with probability at least $1-\delta$, we have:
@@ H.2 Suboptimality Gap Under Sufficient Diversity Condition
$\textup{{SubOpt}}(T)=\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{\lambda_{0}(T-|{\mathcal{T}}^{w}|)K}}\right)=\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{\lambda_{0}\big(T-\min\{(dK)^{2}/\kappa,T-1\}\big)K}}\right).$
Discussion of Theorem H.1. Theorem H.1 shows that for sufficiently large $T$ (i.e., $T=\Omega((dK)^{2}/\kappa+\log(dT)/\lambda_{0})$), the suboptimality gap under the uniform random assortment selection strategy achieves $\widetilde{\mathcal{O}}\big(\sqrt{\frac{d}{\lambda_{0}TK}}\big)$. This result suggests that when the feature space is sufficiently diverse, uniform random selection is effective for learning. It also provides a theoretical explanation for the empirical success of many RLHF implementations [76, 59], where the feature space is sufficiently diverse and prompt-action (sub)set pairs are often selected uniformly at random.
Note that the lower bound we establish in Theorem 3 does not rely on the diversity assumption (Assumption H.1). As a result, deriving a lower bound under the diversity assumption remains an open question, which we leave for future work.
Proof of Theorem H.1.
To provide the proof of Theorem H.1, we first introduce useful concentration inequalities.
Lemma H.1 (Matrix Chernoff, Adapted Sequence from Tropp 80).
Consider a finite adapted sequence $\{X_{k}\}$ with filtratio $\{\mathcal{F}_{t}\}_{t\geq 0}$ of positive-semi definite matrices with dimension $d$, and suppose that $\lambda_{\max}(X_{k})\leq R$ almost surely. Define the finite series
$Y:=\sum_{k}X_{k},\quad\text{and}\quad W:=\sum_{k}\mathbb{E}_{k-1}X_{k}.$
Then, for all $\mu\geq 0$, we have
$\mathbb{P}\left\{\lambda_{\min}(Y)\leq(1-\delta)\mu\,\,\,\text{and}\,\,\,\lambda_{\min}(W)\geq\mu\right\}\leq d\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{\mu/R},\quad\delta\in[0,1).$
By setting $\delta=\frac{3}{4}$, $\mu=\lambda_{0}t$, $R=x_{\max}$ in Lemma H.1, we obtain the following result.
Corollary H.1 (Eigenvalue Growth of Adaptive Gram Matrix).
If $\|X_{i}\|_{2}\leq x_{\max}$ and $\lambda_{\min}\left(\mathbb{E}\left[X_{i}X_{i}^{\top}\mid\mathcal{F}_{i-1}\right]\right)\geq\lambda_{0}$, then, with probability at least $1-d\exp\left(-c_{1}\frac{\lambda_{0}t}{x_{\max}}\right)$,
$\lambda_{\min}\left(\sum_{i=1}^{t}X_{i}X_{i}^{\top}\right)\geq\frac{\lambda_{0}}{4}t$
holds for some absolute constant $c_{1}$.
Now we are ready to provide the proof of Theorem H.1. For simplicity, we present only the case of the PL loss, since the extension to the RB loss directly follows from similar arguments in the proof of Theorem 2. By the definition of the suboptimality gap, we have
@@
$\leq\mathbb{E}_{x\sim\rho}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$ ($\widehat{\pi}_{T}(x)=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi(x,a)^{\top}\widehat{\bm{\theta}}_{T+1}$)
@@
$\leq\left\|\mathbb{E}_{x\sim\rho}\left[\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right]\right\|_{H_{T+1}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ (Hölder’s ineq.)
@@
$\leq\beta_{T+1}(\delta)\left\|\mathbb{E}_{x\sim\rho}\left[\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right]\right\|_{H_{T+1}^{-1}}.$ (Corollary D.1, with prob. $1-\delta$)
To proceed, we slightly modify the definition of $\Lambda_{t}$, as we no longer compute the reference action explicitly. Let $a_{s}^{\prime}$ be an arbitrary action selected from $S_{s}$, which can simply be chosen by sampling uniformly from $S_{s}$. Additionally, the regularization term $\lambda$ is no longer required. Then, we redefine $\Lambda_{t}$ as follows:
$\Lambda_{t}:=\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime}_{s})\big)^{\top},\quad a^{\prime}_{s}\in S$
where
@@
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\right\}.$
Then, by Lemma D.2, we obtain
$\left\|\mathbb{E}_{x\sim\rho}\left[\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right]\right\|_{H_{T+1}^{-1}}$ $\leq\sqrt{50}\left\|\mathbb{E}_{x\sim\rho}\left[\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right]\right\|_{\Lambda_{T+1}^{-1}}$ (Lemma D.2, with prob. $1-\delta$)
$\leq\frac{10\sqrt{2}}{\lambda_{\min}\left(\Lambda_{T+1}\right)}$ ($\|\phi(x,a)\|_{2}\leq 1$)
Under the diversity assumption (Assumption H.1), for $T\geq\frac{c}{\lambda_{0}}\log\frac{d}{\delta}$ with some constant $c>0$, by Corollary H.1, we have, with probability at least $1-\delta$,

### v2

@@ F.1 Main Proof of Theorem 3
Theorem F.1 (Lower bound on adaptive PL model parameter estimation).
@@
Proof of Theorem F.1.
The analysis of this result closely follows the proof of Theorem 3 in shamir2013complexity. The key distinction lies in the input structure: our setting involves a set of feature vectors, while theirs is restricted to a single feature vector.
@@
$\geq\mu^{2}\cdot\mathbb{E}_{\bm{\theta}\sim\operatorname{Unif}(\Theta)}\mathbb{E}_{\bm{\theta}}\left[\sum_{i=1}^{d}\mathbb{I}\left\{\bm{\theta}_{i}\widehat{\bm{\theta}}_{i}<0\right\}\right].$ (F.1)
As in shamir2013complexity, we assume that the query strategy is deterministic conditioned on the past: that is, $S_{t}$ is a deterministic function of the previous queries and observations, i.e., $S_{1},\sigma_{1},\dots,S_{t-1},\sigma_{t-1}$. This assumption is made without loss of generality, since any randomized querying strategy can be viewed as a distribution over deterministic strategies. Therefore, a lower bound that holds uniformly for all deterministic strategies also applies to any randomized strategy. Then, we use the following lemma.
Lemma F.1 (Lemma 4 of shamir2013complexity).
@@
The proof is deferred to Appendix F.2.1.
By applying Lemma F.2, we have
@@
$\geq\frac{d}{2}\left(1-\sqrt{\frac{2K^{2}T\mu^{2}}{d}}\right).$ (F.2)
Combining Equation (F.1) and (F.2), we prove the first inequality of Theorem F.1. The second inequality directly follows by choosing $\mu=\sqrt{d/(8K^{2}T)}$. ∎
@@
The structure of our proof is similar to that of Theorem 2 in wagenmaker2022reward. However, while they consider the linear bandit setting, we focus on the Plackett–Luce (PL) bandit setting.
We adopt the same instance construction as in Theorem F.1, where $\Phi={\mathcal{S}}^{d-1}$ and $\Theta=\{-\mu,\mu\}^{d}$. Define $\phi^{\star}(\bm{\theta})=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi_{a}^{\top}\bm{\theta}$. Then, since $\phi^{\star}(\bm{\theta})\in\Phi$ and $\bm{\theta}\in\Phi$, it is clear that
$\phi^{\star}(\bm{\theta})=\bm{\theta}/\|\bm{\theta}\|_{2}=\bm{\theta}/(\sqrt{d}\mu),\quad\text{and }\,\phi^{\star}(\bm{\theta})^{\top}\bm{\theta}=\sqrt{d}\mu.$ (F.3)
@@
$\mathbb{E}_{\phi\sim\pi}\left[\phi^{\top}\bm{\theta}\right]=\big(\underbrace{\mathbb{E}_{\phi\sim\pi}\left[\phi\right]}_{=:\phi_{\pi}}\big)^{\top}\bm{\theta}\geq\phi^{\star}(\bm{\theta})^{\top}\bm{\theta}-\epsilon=\sqrt{d}\mu-\epsilon.$ (F.4)
@@
$\Longleftrightarrow\,\,$ $\bm{\theta}^{\star}\Delta\leq-\frac{\sqrt{d}\mu}{2}\|\Delta\|_{2}^{2}.$ (Eqn. (F.3))
@@
$-\epsilon\leq-\frac{\sqrt{d}\mu}{2}\|\Delta\|_{2}^{2}.$ (Eqn. (F.4))
@@
$\big\|\widehat{\bm{\theta}}-\bm{\theta}\big\|_{2}=\big\|\sqrt{d}\mu(\widehat{\phi}-\Delta^{\prime})-\sqrt{d}\mu(\widehat{\phi}-\Delta)\big\|_{2}\leq 2\sqrt{d}\mu\sqrt{\frac{2\epsilon}{\sqrt{d}\mu}}=\sqrt{8\sqrt{d}\mu\epsilon}.$ (F.5)
@@
$\leq 8\sqrt{d}\mu\epsilon+\mathbb{E}_{\bm{\theta}}\left[\big\|\widehat{\bm{\theta}}-\bm{\theta}\big\|_{2}^{2}\cdot\mathbb{I}\{\mathcal{E}^{c}\}\right]$ (Eqn. (F.5))
@@
On the other hand, by Theorem F.1, there exists a parameter $\bm{\theta}\in\Theta$ such that, if we collect $T$ samples and set $\mu=\sqrt{d/(8K^{2}T)}$, then the following lower bound holds:
@@
F.2 Proof of Lemmas for Theorem 3
F.2.1 Proof of Lemma F.2
Proof of Lemma F.2.
@@ F.2 Proof of Lemmas for Theorem 3
$D_{\text{KL}}\big(\mathbb{P}_{\bm{\theta}}(\cdot|S)\|\mathbb{P}_{\bm{\theta}^{\prime}}(\cdot|S)\big)=\mathbb{E}_{\sigma\sim\mathbb{P}_{\bm{\theta}}(\cdot|S)}\left[\sum_{j=1}^{|S|}\left(\phi_{\sigma_{j}}^{\top}\left(\bm{\theta}-\bm{\theta}^{\prime}\right)-\log\frac{\sum_{k=j}^{|S|}e^{\phi_{\sigma_{k}}^{\top}\bm{\theta}}}{\sum_{k=j}^{|S|}e^{\phi_{\sigma_{k}}^{\top}\bm{\theta}^{\prime}}}\right)\right].$ (F.6)
@@
$\leq\sum_{k=j}^{|S|}p_{k}(\bm{\theta})\phi_{\sigma_{k}}^{\top}\left(\bm{\theta}^{\prime}-\bm{\theta}\right)+\frac{1}{2}\sum_{a\in S}\left(\phi_{a}^{\top}(\bm{\theta}^{\prime}-\bm{\theta})\right)^{2},$ (F.7)
@@
Plugging Equation (F.7) into Equation (F.6), we get
@@
Appendix G Additional Discussions
This section provides further discussion of our approach. In Subsection G.1, we show how it extends to the regret-minimization setting and improves the existing $\mathcal{O}(e^{B})$ dependence. In Subsection G.2, we describe how the approach also applies to the active-learning setting considered in [das2024active].
G.1 Avoiding $e^{B}$ Scaling in Regret Minimization
In this subsection, we show that our technique for eliminating the $\mathcal{O}(e^{B})$ dependence in the leading term—primarily by dividing the total rounds into warm-up and non–warm-up phases—can also be applied in the regret-minimization setting (e.g., bengs2022stochastic).
First, we formally define the cumulative regret as:
$\textup{{Reg}}_{T}:=\sum_{t=1}^{T}\Big(r_{\bm{\theta}^{\star}}\left(x_{t},\pi^{\star}(x_{t})\right)-\max_{a\in S_{t}}r_{\bm{\theta}^{\star}}(x_{t},a)\Big),$
where the contexts $x_{t}$ can be given arbitrarily (in contrast to the main paper, where they are drawn from a fixed distribution). We assume the linear rewards (Assumption 1) as in the main paper. Moreover, we use the RB update rule (Procedure 2) in this subsection.
Next, we define
$g_{t}(S):=\frac{1}{|S|}\sum_{a\in S}\dot{\mu}\left((\phi(x_{t},a)-\phi(x_{t},\widehat{a}_{t}))^{\top}\widehat{\bm{\theta}}_{t}\right)\left\|\phi(x_{t},a)-\phi(x_{t},\widehat{a}_{t})\right\|_{H_{t}^{-1}}^{2},$
where $\widehat{a}_{t}:=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widehat{\bm{\theta}}_{t}$. The quantity measures the average uncertainty between the feature vectors of actions in $S$ and that of the current best action $\widehat{a}_{t}$.
We then select the assortment $S_{t}$ in a greedy manner, following the procedure in Equation (11). We initialize $S=\{\widehat{a}_{t}\}$, and subsequently add actions one by one. Specifically, at each step, we select
$a^{\star}\in\mathop{\mathrm{argmax}}_{a\in\mathcal{A}\setminus S}\ \Delta_{t}(a\mid S),\qquad\text{where }\,\Delta_{t}(a\mid S):=g_{t}(S\cup\{a\})-g_{t}(S),$ (G.1)
and include $a^{\star}$ in $S$ if $\Delta_{t}(a\mid S)\geq 0$. The process continues until either $|S|=K$ or no action yields a non-negative gain. Since $g_{t}(\{\widehat{a}_{t}\})=0$, this rule guarantees that the selected set satisfies $|S|\geq 2$.
Theorem G.1 (Regret upper bound).
In the same setting as Theorem 2, suppose $T\geq e^{K/d}$. Then, with probability at least $1-\delta$, the assortment selection rule in Equation (G.1), combined with the RB loss update strategy (Procedure 2), guarantees
@@ G.1 Avoiding $e^{B}$ Scaling in Regret Minimization
$\textup{{Reg}}_{T}=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}}{\kappa}\right).$
Proof of Theorem G.1.
By the definition of the cumulative regret, we have
$\textup{{Reg}}_{T}$ $:=\sum_{t=1}^{T}\left(\phi(x_{t},\pi^{\star}(x_{t}))-\max_{a\in S_{t}}\phi(x_{t},a)\right)^{\top}\bm{\theta}^{\star}$
$\leq\sum_{t=1}^{T}\left(\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\right)^{\top}\bm{\theta}^{\star}$ ($\widehat{a}_{t}\in S_{t}$)
@@
$\leq\sum_{t=1}^{T}\left(\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\right)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t}\right).$ ( $\widehat{a}_{t}:=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widehat{\bm{\theta}}_{t}$)
Recall the definition of the set of warm-up rounds, i.e., ${\mathcal{T}}^{w}=\big\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq 1/\beta_{T}(\delta)\big\}.$ Then, we get
@@
$\textup{{Reg}}_{T}$ $\leq 4B|{\mathcal{T}}^{w}|+\sum_{t\notin{\mathcal{T}}^{w}}\left(\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\right)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t}\right)$
$\leq\mathcal{O}\left(\frac{B}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{T}{d\lambda}\right)\right)+\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\right\|_{H_{t}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t}\right\|_{H_{t}}$ (Lemma D.2 and Cauchy-Schwartz ineq.)
@@
$\leq\mathcal{O}\left(\frac{B}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{T}{d\lambda}\right)\right)+\beta_{T}(\delta)\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\right\|_{H_{t}^{-1}}.$ (Corollary D.1)
For simplify the presentation, let $\psi_{t}(a,a^{\prime})=\phi(x_{t},a)-\phi(x_{t},a^{\prime})$. Then, for $t\notin{\mathcal{T}}^{w}$, we have
$\frac{1}{\dot{\mu}\left(\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\bm{\theta}^{\star}\right)}$ $=\frac{\left(1+e^{\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\bm{\theta}^{\star}}\right)^{2}}{e^{\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\bm{\theta}^{\star}}}$
$\leq\left(1+e^{\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\bm{\theta}^{\star}}\right)^{2}$ ($\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\bm{\theta}^{\star}\geq 0$)
$\leq\left(1+e^{\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t})}\right)^{2}$ ( $\widehat{a}_{t}:=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widehat{\bm{\theta}}_{t}$)
$\leq\left(1+e^{\|\psi_{t}(\pi^{\star}(x_{t}),\widehat{a}_{t})^{\top}\|_{H_{t}^{-1}}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{t}\|_{H_{t}}}\right)^{2}$
@@
$\leq\left(1+e^{\frac{\beta_{t}(\delta)}{\beta_{T}(\delta)}}\right)^{2}\leq(1+e)^{2}.$ ($t\notin{\mathcal{T}}^{w}$ and Corollary D.1)
Therefore, the same line of reasoning used in the proof of Theorem 2 applies here as well.
$\sum_{t\notin{\mathcal{T}}^{w}}$ $\|\phi(x_{t},\pi^{\star}(x_{t}))-\phi(x_{t},\widehat{a}_{t})\|_{H_{t}^{-1}}$